# for functionals in the $\mathcal{A}$-free setting

### v1

@@ 3. $\Gamma$-convergence in the unconstrained setting
Let $u\colon D\to\mathbb{R}^{d}$ be a piecewise constant function, that is, there exists a finite family $(B_{i})_{i\in I}$ of pairwise disjoint sets in $\mathcal{O}(D)$, covering almost all of $D$, and a finite family $(\xi_{i})_{i\in I}$ in $\mathbb{R}^{d}$ such that for every $i\in I$ we have $u(x)=\xi_{i}$ for every $x\in B_{i}$. By (2.10) and (2.12), the set function $B\mapsto F(u,B)$ is a measure that is absolutely continuous with respect to the Lebesgue measure. By applying (3.25) to $\xi_{i}$ and $B_{i}$, we obtain
@@
for every $k\in\mathbb{N}$. By (2.6), (2.8), (2.10), and (2.14), we can pass to the limit in both terms as $k\to\infty$ and we obtain (3.24).
If $F\in\mathcal{I}_{\mathrm{sc}}$, then for a.e. $x\in\mathbb{R}^{N}$ the function $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex by [19, Theorem 3.6]. Fix $x\in\mathbb{R}^{N}$ and $\rho>0$, and let $g\colon\mathbb{R}^{N}\to\mathbb{R}$ be the function defined by
@@
Hence, $g$ is $\mathcal{A}$-quasiconvex which, by (3.24), is equivalent to saying that the function
@@
is $\mathcal{A}$-quasiconvex. Moreover, by Fatou’s lemma for bounded sequences (see [17, Lemma 1.83 (ii)]), it can be similarly checked that the $\limsup$ of locally equi-bounded functions preserves $\mathcal{A}$-quasiconvexity. We then deduce from (3.23) that $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for every $x\in\mathbb{R}^{N}$. ∎
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}$ and let $(F_{k})_{k\in\mathbb{N}}$ be the corresponding sequence of functionals in $\mathcal{I}$ defined by (2.13). Then, there exist a subsequence, which we do not relabel, and a function $f\in\mathcal{F}_{\mathrm{qc}}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F(\cdot,D)$ defined by (2.13) with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$.
@@
The result follows from Theorems 3.1 and $3.$ ∎
@@ 4. $\Gamma$-convergence in the $\mathcal{A}$-free setting
In this section, we study $\Gamma$-convergence in the $\mathcal{A}$-free setting, i.e., with the constraint $\mathcal{A}u=0$. We begin with some preliminary lemmas. The following result has been established in [19, Lemma 2.15].
@@
$u_{k}\rightharpoonup u\text{ weakly in $L^{p}(D;\mathbb{R}^{d})$}\enspace\text{ and }\enspace\mathcal{A}_{D}u_{k}\to 0\text{ in $W^{-1,p}(D;\mathbb{R}^{l})$.}$
@@
$\begin{gathered}v_{k}\rightharpoonup u\text{ weakly in $L^{p}(D;\mathbb{R}^{d})$},\quad\mathcal{A}_{D}v_{k}=0,\quad\int_{D}v_{k}\,dx=\int_{D}u\,dx,\\ \lim_{k\to\infty}\|u_{k}-v_{k}\|_{L^{r}(D;\mathbb{R}^{d})}=0\enspace\text{ for all $1\leq r<p$}.\end{gathered}$ (4.1)
@@
Let $D$, $u$, $(u_{k})_{k\in\mathbb{N}}$, and $(v_{k})_{k\in\mathbb{N}}$ be as in Lemma 4.1, and let $(F_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{I}$. Then,
@@
$\liminf_{k\to\infty}F_{k}(v_{k},D)\leq\liminf_{k\to\infty}F_{k}(u_{k},D)\enspace\hbox{ and }\enspace\limsup_{k\to\infty}F_{k}(v_{k},D)\leq\limsup_{k\to\infty}F_{k}(u_{k},D).$ (4.3)
@@
This result was established in [7] when either $F_{k}$ are independent of $k$ or $F_{k}$ are the functionals associated to functions $f_{k}$ with $f_{k}(x,\xi)=f(kx,\xi)$ for some $f\in\mathcal{F}$ periodic in the first variable. To prove that (4.2) also holds in our setting, we apply Theorem 3.3 to obtain a sequence $(f_{k})_{k\in\mathbb{N}}$ in $\mathcal{F}$ such that
@@
The equalities $\mathcal{A}v_{\rho}=0$ in $\mathbb{R}^{N}$ and $\int_{Q_{\rho}}v_{\rho}\,dx=\int_{Q_{\rho}}u_{\rho}\,dx$ can be obtained from the corresponding properties of $v$ and $u$ by a change of variables. As for (4.6) for $v_{\rho}$ and $u_{\rho}$, given $F_{\rho}\in\mathcal{I}$, let $f_{\rho}\in\mathcal{F}$ be the corresponding integrand (see Theorem 3.3). Let $f\in\mathcal{F}$ be the function defined by
@@
and denote by $F\in\mathcal{I}$ the corresponding functional (see Remark 2.8). Then, by (4.6) for $Q$, we have
@@
Since $F_{\rho}$ is an arbitrary element of $\mathcal{I}$, we obtain (4.6) for $v_{\rho}$ and $u_{\rho}$. This concludes the proof of the claim that it is enough to prove the corollary for $Q:=Q_{1}(0)$, which we establish next.
We argue by contradiction. Assume that the statement for $Q$ is false. Then, there exist $\varepsilon>0$ and $C>0$ such that for every $k\in\mathbb{N}$ there exists $u_{k}\in L^{p}(Q;\mathbb{R}^{d})$, with $\|u_{k}\|_{L^{p}(Q;\mathbb{R}^{d})}^{p}<C$, $\operatorname{supp}u_{k}\subset\subset Q$, and $\|\widetilde{\mathcal{A}}_{Q}u_{k}\|_{\widetilde{W}^{-1,p}(Q;\mathbb{R}^{l})}^{p}<1/k$, such that for every $v\in L^{p}_{\rm per}(Q;\mathbb{R}^{d})$, with $\|v-u_{k}\|_{W^{-1,p}(Q;\mathbb{R}^{d})}^{p}<\varepsilon$, $\mathcal{A}v=0$ in $\mathbb{R}^{N}$, and $\int_{Q}v\,dx=\int_{Q}u_{k}\,dx$, there exists $F_{k,v}\in\mathcal{I}$ such that
@@
Let $(F_{k})_{k\in\mathbb{N}}$ be a sequence of functionals in $\mathcal{I}$, let $F\in\mathcal{I}$, and let $D\in\mathcal{O}(\mathbb{R}^{N})$. If we extend $F^{\mathcal{A}}(\cdot,D)$ to $L^{p}(D;\mathbb{R}^{d})$ by setting $F^{\mathcal{A}}(u,D)=+\infty$ for every $u\in L^{p}(D;\mathbb{R}^{d})\setminus\ker\mathcal{A}_{D}$, by (2.10) we can apply [11, Proposition 8.16] to conclude that the sequence $(F_{k}^{\mathcal{A}}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F^{\mathcal{A}}(\cdot,D)$ in $\ker\mathcal{A}_{D}$ with respect to the weak topology of $L^{p}(D;\mathbb{R}^{d})$ if and only if
@@
Let $(F_{k})_{k\in\mathbb{N}}$ be a sequence of functionals in $\mathcal{I}$, let $F\in\mathcal{I}$, and let $F_{k}^{\mathcal{A}}$ and $F^{\mathcal{A}}$ be as in Definition 4.4. The following conditions are equivalent:
@@
which exists by [11, Proposition 8.1]. We observe further that $u_{k}\rightharpoonup u$ weakly in $L^{p}(D;\mathbb{R}^{d})$ (see Remark 2.7). Let $(v_{k})_{k\in\mathbb{N}}$ be the sequence in $\ker\mathcal{A}_{D}$ provided by Lemma 4.1. By Lemma 4.2, we have
@@
Assume (b). By the Theorem 3.1, there exist a subsequence, which we do not relabel, and a functional $\hat{F}\in\mathcal{I}_{\mathrm{sc}}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $\hat{F}(\cdot,D)$ with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$. Since (a) implies (b), we have $\hat{F}(u,D)=F(u,D)$ for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u\in\ker\mathcal{A}_{D}$. Let $\hat{f}$ and $f$ be the integrands corresponding to $\hat{F}$ and $F$, respectively, defined by (3.23). Since every constant function belongs to $\ker\mathcal{A}_{D}$, we deduce that $\hat{f}=f$; by Theorem 3.3, this implies $\hat{F}=F$. Since the $\Gamma$-limit does not depend on the subsequence, we obtain (a) by the Urysohn property of $\Gamma$-convergence (see [11, Proposition 8.3]). ∎
@@
Let $(F_{k})_{k\in\mathbb{N}}$ be a sequence of functionals in $\mathcal{I}$ and let $F_{k}^{\mathcal{A}}$ be as in Definition 4.4. Then, there exist a subsequence, which we do not relabel, and a functional $F\in\mathcal{I}_{\mathrm{sc}}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}^{\mathcal{A}}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F^{\mathcal{A}}(\cdot,D)$ in $\ker\mathcal{A}_{D}$ with respect to the weak topology of $L^{p}(D;\mathbb{R}^{d})$.

### v2

@@ 3. $\Gamma$-convergence in the unconstrained setting
Let $u\colon D\to\mathbb{R}^{d}$ be a piecewise constant function, that is, there exists a finite family $(B_{i})_{i\in I}$ of pairwise disjoint sets in $\mathcal{O}(D)$, covering almost all of $D$, and a finite family $(\xi_{i})_{i\in I}$ in $\mathbb{R}^{d}$ such that for every $i\in I$ we have $u(x)=\xi_{i}$ for every $x\in B_{i}$. By conditions (a) and (b) of Definition 2.9 the set function $B\mapsto F(u,B)$ is a measure that is absolutely continuous with respect to the Lebesgue measure. By applying (3.25) to $\xi_{i}$ and $B_{i}$, we obtain
@@
for every $k\in\mathbb{N}$. By condition (a) of Definition 2.9, using (2.10), (2.13), and (2.15) we can pass to the limit in both terms as $k\to\infty$ and we obtain (3.24).
If $F\in\mathcal{I}_{\mathrm{sc}}$, then for a.e. $x\in\mathbb{R}^{N}$ the function $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex by [20, Theorem 3.6]. Fix $x\in\mathbb{R}^{N}$ and $\rho>0$, and let $g\colon\mathbb{R}^{N}\to\mathbb{R}$ be the function defined by
@@
Hence, $g$ is $\mathcal{A}$-quasiconvex. By (3.24) this is equivalent to saying that the function
@@
is $\mathcal{A}$-quasiconvex. Moreover, by Fatou’s lemma for bounded sequences (see [18, Lemma 1.83 (ii)]), it can be similarly checked that the $\limsup$ of locally equi-bounded functions preserves $\mathcal{A}$-quasiconvexity. We then deduce from (3.23) that $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for every $x\in\mathbb{R}^{N}$. ∎
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}_{\mathrm{Lip}}$ and let $(F_{k})_{k\in\mathbb{N}}$ be the corresponding sequence of functionals in $\mathcal{I}$ defined by (2.14). Then, there exist a subsequence, which we do not relabel, and a function $f\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F(\cdot,D)$ defined by (2.14) with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$.
@@
The result follows from Remark 2.11 and Theorems 3.1 and $3.$ ∎
The following theorem, which was communicated to us by Jean-François Babadjian, shows that the hypothesis of $p$-Lipschitz continuity can be omitted under a condition on the wave cone $\Lambda$ defined by (2.6).
Theorem 3.5.
Assume that the vector space $\mathrm{span}(\Lambda)$ generated by the wave cone $\Lambda$ coincides with $\mathbb{R}^{d}$. Then, Theorem 3.3 is still satisfied if we replace $\mathcal{I}_{\mathrm{Lip}}$ by $\mathcal{I}_{\mathrm{sc}}$ and $\mathcal{F}_{\mathrm{Lip}}$ by $\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$. Moreover, the inclusion $\mathcal{I}_{\mathrm{sc}}\subset\mathcal{I}_{\mathrm{Lip}}$ holds. Finally, Corollary 3.4 remains valid if we assume only that $(f_{k})_{k\in\mathbb{N}}$ is a sequence in $\mathcal{F}$.
Proof.
Assume that $F\in\mathcal{I}_{\mathrm{sc}}$ and let $f$ be defined by (3.23). Arguing as in Theorem 3.3, we show that the function $x\mapsto f(x,\xi)$ is measurable on $\mathbb{R}^{N}$ for every $\xi\in\mathbb{R}^{d}$, that $f$ satisfies the $p$-growth condition (2.10), and that (3.24) holds for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every piecewise constant function $u\colon D\to\mathbb{R}^{d}$.
We claim that for every $x\in\mathbb{R}^{N}$ the function $\xi\mapsto f(x,\xi)$ is $\Lambda$-convex on $\mathbb{R}^{d}$, i.e.,
$f(x,\theta\xi_{1}+(1-\theta)\xi_{2})\leq\theta f(x,\xi_{1})+(1-\theta)f(x,\xi_{2})$ (3.27)
for every $\xi_{1}$, $\xi_{2}\in\mathbb{R}^{d}$ with $\xi_{1}-\xi_{2}\in\Lambda$ and every $\theta\in[0,1]$. To this aim, we fix $\xi_{1}$, $\xi_{2}$, and $\theta$ as required and we consider the periodic extension $\chi\colon\mathbb{R}\to[0,1]$ of the characteristic function of $0,\theta]$. By the definition of wave cone (see (2.6)), there exists $w\in\mathbb{R}^{N}\setminus\{0\}$ such that
$\xi_{1}-\xi_{2}\in\ker\left(\sum_{i=1}^{N}A^{i}w_{i}\right).$ (3.28)
We consider the piecewise constant function $u_{\varepsilon}\in L^{p}_{\mathrm{loc}}(\mathbb{R}^{N};\mathbb{R}^{d})$ defined by
$u_{\varepsilon}(x)=\chi(x\cdot w/\varepsilon)\xi_{1}+(1-\chi(x\cdot w/\varepsilon))\xi_{2}.$
By the Riemann–Lebesgue Lemma, $u_{\varepsilon}\rightharpoonup\theta\xi_{1}+(1-\theta)\xi_{2}$ weakly in $L^{p}(D;\mathbb{R}^{d})$ for every $D\in\mathcal{O}(\mathbb{R}^{N})$. Moreover,
[…]
Since $F\in\mathcal{I}_{\mathrm{sc}}$, for every $D\in\mathcal{O}(\mathbb{R}^{N})$ the functional $F(\cdot,D)$ is lower semicontinuous with respect to the norm $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$. Thus, by Remark 2.10,
$F(\theta\xi_{1}+(1-\theta)\xi_{2}),D)\leq\liminf_{\varepsilon\to 0+}F(u_{\varepsilon},D).$
Using the integral representation of $F$ on piecewise constant functions, from the previous inequality we obtain
$F(\theta\xi_{1}+(1-\theta)\xi_{2}),D)$ $\leq\liminf_{\varepsilon\to 0+}\int_{D}f(x,u_{\varepsilon}(x))\,dx$
$=\lim_{\varepsilon\to 0+}\int_{D}\big(\chi(x\cdot w/\varepsilon)f(x,\xi_{1})+(1-\chi(x\cdot w/\varepsilon))f(x,\xi_{2})\big)\,dx$
$=\theta\int_{D}f(x,\xi_{1})\,dx+(1-\theta)\int_{D}f(x,\xi_{2})\,dx=\theta F(\xi_{1},D)+(1-\theta)F(\xi_{2},D),$
where the second equality follows again from the Riemann–Lebesgue Lemma.
Given $x\in\mathbb{R}^{N}$ and $\rho>0$, we take $D=Q_{\rho}(x)$ in the previous inequality. Dividing by $\rho^{N}$ and using (3.23), we obtain (3.27), which shows that the function $\xi\mapsto f(x,\xi)$ is $\Lambda$-convex on $\mathbb{R}^{d}$ for every $x\in\mathbb{R}^{N}$.
By (2.10), we can apply [23, Lemma 2.3]) (see also [21, Lemma 4.6]) to obtain that there exists $\check{c}_{1}>0$, depending only on $c_{0}$, $p$, and $d$, such that (2.13) holds for every $x\in\mathbb{R}^{N}$ and every $\xi_{1},\xi_{2}\in\mathbb{R}^{d}$. Since $f$ is measurable with respect to $x$ and satisfies the growth conditions (2.10), we conclude that $f\in\mathcal{F}$.
Thanks to (2.13), the integral representation on piecewise constant functions leads to the following inequality
$F(u,D)\leq\int_{D}f(x,u(x))\,dx\quad\hbox{for all }D\in\mathcal{O}(\mathbb{R}^{N})\hbox{ and }u\in L^{p}(D;\mathbb{R}^{d})$
by the lower semicontinuity of $F(\cdot,D)$ in the strong topology of $L^{p}(D;\mathbb{R}^{d})$ (see Remark 2.10) and the continuity of $u\mapsto\int_{D}f(x,u(x))\,dx$ in the strong topology of $L^{p}(D;\mathbb{R}^{d})$ due to (2.13).
To prove the equality, we fix $u\in L^{p}_{\mathrm{loc}}(\mathbb{R}^{N};\mathbb{R}^{d})$ and we use the translation argument introduced in the proof of [8, Lemma 4.1] (see also [12, Theorem 21.1]). We define
$G(v,D)=F(u+v,D)\quad\hbox{for all }D\in\mathcal{O}(\mathbb{R}^{N})\hbox{ and }v\in L^{p}(D;\mathbb{R}^{d}).$
We observe that $G$ satisfies the growth conditions
  • (a)
    $\frac{1}{C}\|v\|_{L^{p}(D;\mathbb{R}^{d})}^{p}-C\big(|D|+\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p}\big)\leq G(v,D)\leq C\big(|D|+\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p}\big)+C\|v\|_{L^{p}(D;\mathbb{R}^{d})}^{p},$
for suitable a constant $C>1$. This is a slight variant of condition (a) of Definition 2.9.
Arguing as before, we obtain a Carathéodory function $g\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}$ with $p$-growth and $p$-Lipschitz with respect to the second variable such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$ we have
$G(v,D)=\int_{D}g(x,v(x))\,dx\quad\hbox{if }v\hbox{ is piecewise constant on }D,$
from which it follows that
$G(v,D)\leq\int_{D}g(x,v(x))\,dx\quad\hbox{for all }v\in L^{p}(D;\mathbb{R}^{d}).$
Let $D\in\mathcal{O}(\mathbb{R}^{N})$ and let $(u_{k})_{k\in\mathbb{N}}$ be a sequence of piecewise constant functions such that $u_{k}\to u$ strongly in $L^{p}(D;\mathbb{R}^{d})$. Since $v\mapsto\int_{D}f(x,v(x))\,dx$ and $v\mapsto\int_{D}g(x,v(x))\,dx$ are continuous in the strong topology of $L^{p}(D;\mathbb{R}^{d})$, the equalities and inequalities satisfied by $F(v,D)$, $G(v,D)$, $\int_{D}f(x,v(x))\,dx$ and $\int_{D}g(x,v(x))\,dx$ give
$F(u,D)$ $\leq\int_{D}f(x,u(x))\,dx=\lim_{k\to\infty}\int_{D}f(x,u_{k}(x))\,dx=\lim_{k\to\infty}F(u_{k},D)=\lim_{k\to\infty}G(u_{k}-u,D)$
$\leq\lim_{k\to\infty}\int_{D}g(x,u_{k}(x)-u(x))\,dx=\int_{D}g(x,0)\,dx=G(0,D)=F(u,D),$
and we obtain
$F(u,D)=\int_{D}f(x,u(x))\,dx.$ (3.30)
Since every $u\in L^{p}(D;\mathbb{R}^{d})$ can be extended to a function of $L^{p}_{\mathrm{loc}}(D;\mathbb{R}^{d})$, (3.24) holds for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and $u\in L^{p}(D;\mathbb{R}^{d})$.
Since $f\in\mathcal{F}$ and $F\in\mathcal{I}_{\mathrm{sc}}$, using [20, Theorem 3.6] we deduce from (3.30) that the function $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for a.e. $x\in\mathbb{R}^{N}$. Arguing as in the last part of the proof of Theorem 3.3, we obtain that $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for every $x\in\mathbb{R}^{N}$, hence $f\in\mathcal{F}_{\mathrm{qc}}$. By Remark 2.8 ,we have $\mathcal{F}_{\mathrm{qc}}\subset\mathcal{F}_{\mathrm{Lip}}$, hence $f\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$. This concludes the proof of the modified version of Theorem 3.3.
To prove the inclusion $\mathcal{I}_{\mathrm{sc}}\subset\mathcal{I}_{\mathrm{Lip}}$, we observe that, by the modified version of Theorem 3.3 for every $F\in\mathcal{I}_{\mathrm{sc}}$ there exists $f\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ such that (3.24) holds for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u\in L^{p}(D,\mathbb{R}^{d})$. By Remark 2.11, this implies that $F\in\mathcal{I}_{\mathrm{Lip}}$.
The modified version of Corollary 3.4 follows easily from this modified version of Theorem 3.3. ∎
@@ 4. $\Gamma$-convergence in the $\mathcal{A}$-free setting
In this section, we study $\Gamma$-convergence in the $\mathcal{A}$-free setting, i.e., with the constraint $\mathcal{A}u=0$. We begin with some preliminary lemmas. The following result has been established in [20, Lemma 2.15].
@@
$u_{k}\rightharpoonup u\text{ weakly in $L^{p}(D;\mathbb{R}^{d})$}\kern 5.0pt\text{ and }\kern 5.0pt\mathcal{A}_{D}u_{k}\to 0\text{ in $W^{-1,p}(D;\mathbb{R}^{l})$.}$
@@
$\begin{gathered}v_{k}\rightharpoonup u\text{ weakly in $L^{p}(D;\mathbb{R}^{d})$},\quad\mathcal{A}_{D}v_{k}=0,\quad\int_{D}v_{k}\,dx=\int_{D}u\,dx,\\ \lim_{k\to\infty}\|u_{k}-v_{k}\|_{L^{r}(D;\mathbb{R}^{d})}=0\kern 5.0pt\text{ for all $1\leq r<p$}.\end{gathered}$ (4.1)
@@
Let $D$, $u$, $(u_{k})_{k\in\mathbb{N}}$, and $(v_{k})_{k\in\mathbb{N}}$ be as in Lemma 4.1, and let $(F_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{I}_{\mathrm{Lip}}$. Then,
@@
$\liminf_{k\to\infty}F_{k}(v_{k},D)\leq\liminf_{k\to\infty}F_{k}(u_{k},D)\kern 5.0pt\hbox{ and }\kern 5.0pt\limsup_{k\to\infty}F_{k}(v_{k},D)\leq\limsup_{k\to\infty}F_{k}(u_{k},D).$ (4.3)
@@
This result was established in [7] when either $F_{k}$ are independent of $k$ or $F_{k}$ are the functionals associated to functions $f_{k}$ with $f_{k}(x,\xi)=f(kx,\xi)$ for some $f\in\mathcal{F}$ periodic in the first variable. To prove that (4.2) also holds in our setting, we apply Theorem 3.3 to obtain a sequence $(f_{k})_{k\in\mathbb{N}}$ in $\mathcal{F}_{\mathrm{Lip}}$ such that
@@
The equalities $\mathcal{A}v_{\rho}=0$ in $\mathbb{R}^{N}$ and $\int_{Q_{\rho}}v_{\rho}\,dx=\int_{Q_{\rho}}u_{\rho}\,dx$ can be obtained from the corresponding properties of $v$ and $u$ by a change of variables. As for (4.6) for $v_{\rho}$ and $u_{\rho}$, given $F_{\rho}\in\mathcal{I}_{\mathrm{Lip}}$, let $f_{\rho}\in\mathcal{F}_{\mathrm{Lip}}$ be the corresponding integrand (see Theorem 3.3). Let $f\in\mathcal{F}_{\mathrm{Lip}}$ be the function defined by
@@
and denote by $F\in\mathcal{I}_{\mathrm{Lip}}$ the corresponding functional (see Remark 2.11). Then, by (4.6) for $Q$, we have
@@
Since $F_{\rho}$ is an arbitrary element of $\mathcal{I}_{\mathrm{Lip}}$, we obtain (4.6) for $v_{\rho}$ and $u_{\rho}$. This concludes the proof of the claim that it is enough to prove the corollary for $Q:=Q_{1}(0)$, which we establish next.
We argue by contradiction. Assume that the statement for $Q$ is false. Then, there exist $\varepsilon>0$ and $C>0$ such that for every $k\in\mathbb{N}$ there exists $u_{k}\in L^{p}(Q;\mathbb{R}^{d})$, with $\|u_{k}\|_{L^{p}(Q;\mathbb{R}^{d})}^{p}<C$, $\operatorname{supp}u_{k}\subset\subset Q$, and $\|\widetilde{\mathcal{A}}_{Q}u_{k}\|_{\widetilde{W}^{-1,p}(Q;\mathbb{R}^{l})}^{p}<1/k$, such that for every $v\in L^{p}_{\rm per}(Q;\mathbb{R}^{d})$, with $\|v-u_{k}\|_{W^{-1,p}(Q;\mathbb{R}^{d})}^{p}<\varepsilon$, $\mathcal{A}v=0$ in $\mathbb{R}^{N}$, and $\int_{Q}v\,dx=\int_{Q}u_{k}\,dx$, there exists $F_{k,v}\in\mathcal{I}_{\mathrm{Lip}}$ such that
@@
Let $(F_{k})_{k\in\mathbb{N}}$ be a sequence of functionals in $\mathcal{I}$, let $F\in\mathcal{I}$, and let $D\in\mathcal{O}(\mathbb{R}^{N})$. If we extend $F^{\mathcal{A}}(\cdot,D)$ to $L^{p}(D;\mathbb{R}^{d})$ by setting $F^{\mathcal{A}}(u,D)=+\infty$ for every $u\in L^{p}(D;\mathbb{R}^{d})\setminus\ker\mathcal{A}_{D}$, by the lower bound in (a) of Definition 2.9 we can apply [12, Proposition 8.16] to conclude that the sequence $(F_{k}^{\mathcal{A}}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F^{\mathcal{A}}(\cdot,D)$ in $\ker\mathcal{A}_{D}$ with respect to the weak topology of $L^{p}(D;\mathbb{R}^{d})$ if and only if
@@
Let $(F_{k})_{k\in\mathbb{N}}$ be a sequence of functionals in $\mathcal{I}_{\mathrm{Lip}}$, let $F\in\mathcal{I}$, and let $F_{k}^{\mathcal{A}}$ and $F^{\mathcal{A}}$ be as in Definition 4.4. The the following conditions are equivalent:
@@
which exists by [12, Proposition 8.1]. We observe further that $u_{k}\rightharpoonup u$ weakly in $L^{p}(D;\mathbb{R}^{d})$ (see Remark 2.10). Let $(v_{k})_{k\in\mathbb{N}}$ be the sequence in $\ker\mathcal{A}_{D}$ provided by Lemma 4.1. By Lemma 4.2, we have
@@
Assume (b). By the Theorem 3.1, there exist a subsequence, which we do not relabel, and a functional $\hat{F}\in\mathcal{I}_{\mathrm{sc}}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $\hat{F}(\cdot,D)$ with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$. Since (a) implies (b), we have $\hat{F}(u,D)=F(u,D)$ for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u\in\ker\mathcal{A}_{D}$. Let $\hat{f}$ and $f$ be the integrands corresponding to $\hat{F}$ and $F$, respectively, defined by (3.23). Since every constant function belongs to $\ker\mathcal{A}_{D}$, we deduce that $\hat{f}=f$; by Theorems 3.3 and 3.5 this implies $\hat{F}=F$. Since the $\Gamma$-limit does not depend on the subsequence, we obtain (a) by the Urysohn property of $\Gamma$-convergence (see [12, Proposition 8.3]). ∎
@@
Let $(F_{k})_{k\in\mathbb{N}}$ be a sequence of functionals in $\mathcal{I}_{\mathrm{Lip}}$ and let $F_{k}^{\mathcal{A}}$ be as in Definition 4.4. Then, there exist a subsequence, which we do not relabel, and a functional $F\in\mathcal{I}_{\mathrm{sc}}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}^{\mathcal{A}}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F^{\mathcal{A}}(\cdot,D)$ in $\ker\mathcal{A}_{D}$ with respect to the weak topology of $L^{p}(D;\mathbb{R}^{d})$.